OpenStudy (anonymous):

find a function whose graph is the given curve: the bottom half of a circle x^2+y^2=16

OpenStudy (perl):

if you solve that for y, what do you get

OpenStudy (perl):

x^2+y^2=16 y^2 = 16 - x^2 y = +/- sqrt( 16 - x^2 ) now you have two graphs here y = sqrt(16 - x^2) y = - sqrt(16 -x^2) determine which one is the bottom half